# Context-free graphs and their transition groups

### v2

@@ 5. Detecting torsion
We give some structural results focusing on the existence of torsion elements and on bounds on the order of such elements. Henceforth, we consider a context-free $A$-digraph $\Gamma$ with $N$ end-cone types $\Gamma_{1},\ldots,\Gamma_{N}$ with the corresponding frontier vertices $\Delta_{1},\ldots,\Delta_{N}$ with $\Delta_{i}\subseteq\Gamma_{i}$. We say that two end-cones $\Gamma(p,x_{0}),\Gamma(q,x_{0})$ are self-similar if $\Gamma(p,x_{0})\subseteq\Gamma(q,x_{0})$ and there is an end-isomorphism $\psi:\Gamma(q,x_{0})\to\Gamma(p,x_{0})$ with $\psi(q)=p$. We have the following lemma.
@@
Let $\Gamma(p,x_{0}),\Gamma(q,x_{0})$ be two distinct end-isomorphic self-similar end-cones of $\Gamma$. Then, for any walk $q\smash{\stackrel{{\scriptstyle\text{\tiny{$u$}}}}{{\relbar\joinrel\longrightarrow}}}p\subseteq\Gamma(q,x_{0})$ there is an infinite walk $q\smash{\stackrel{{\scriptstyle\text{\tiny{$u^{\omega}$}}}}{{\relbar\joinrel\longrightarrow}}}\subseteq\Gamma(q,x_{0})$ that is not a circuit.
@@
Indeed, let $\psi:\Gamma(q,x_{0})\to\Gamma(p,x_{0})$ be the end-isomorphism with $\psi(q)=p$. Consider a geodesic $q\smash{\stackrel{{\scriptstyle\text{\tiny{$g$}}}}{{\relbar\joinrel\longrightarrow}}}\psi(q)\subseteq\Gamma(q,x_{0})$ connecting $q$ with $\psi(q)=p$. We claim that also
$q\smash{\stackrel{{\scriptstyle\text{\tiny{$g$}}}}{{\relbar\joinrel\longrightarrow}}}\psi(q)\smash{\stackrel{{\scriptstyle\text{\tiny{$g$}}}}{{\relbar\joinrel\longrightarrow}}}\psi^{2}(q)$
@@
is a geodesic. Any geodesic $q\smash{\stackrel{{\scriptstyle\text{\tiny{$u$}}}}{{\relbar\joinrel\longrightarrow}}}\psi^{2}(q)\subseteq\Gamma(q,x_{0})$ has to cross a vertex $v\in\Delta(\psi(q),x_{0})$, thus the previous geodesic factors as $q\smash{\stackrel{{\scriptstyle\text{\tiny{$u_{1}$}}}}{{\relbar\joinrel\longrightarrow}}}v\smash{\stackrel{{\scriptstyle\text{\tiny{$u_{2}$}}}}{{\relbar\joinrel\longrightarrow}}}\psi^{2}(q)$ with $q\smash{\stackrel{{\scriptstyle\text{\tiny{$u_{1}$}}}}{{\relbar\joinrel\longrightarrow}}}v\subseteq\Gamma(q,x_{0})$ and $v\smash{\stackrel{{\scriptstyle\text{\tiny{$u_{2}$}}}}{{\relbar\joinrel\longrightarrow}}}\psi^{2}(q)\subseteq\Gamma(\psi(q),x_{0})$. Since both $q\smash{\stackrel{{\scriptstyle\text{\tiny{$g$}}}}{{\relbar\joinrel\longrightarrow}}}\psi(q)$ and $\psi(q)\smash{\stackrel{{\scriptstyle\text{\tiny{$g$}}}}{{\relbar\joinrel\longrightarrow}}}\psi^{2}(q)$ are geodesics, respectively in $\Gamma(q,x_{0})$ and $\Gamma(\psi(q),x_{0})$ we deduce $|u_{1}|=|g|$ and $|u_{2}|=|g|$. Hence, $q\smash{\stackrel{{\scriptstyle\text{\tiny{$g$}}}}{{\relbar\joinrel\longrightarrow}}}\psi(q)\smash{\stackrel{{\scriptstyle\text{\tiny{$g$}}}}{{\relbar\joinrel\longrightarrow}}}\psi^{2}(q)$ is also a geodesic. The same argument can be extended to show that $q\smash{\stackrel{{\scriptstyle\text{\tiny{$g^{k}$}}}}{{\relbar\joinrel\longrightarrow}}}\psi^{k}(q)\subseteq\Gamma(q,x_{0})$ is a geodesic for any $k\geq 1$. Now, by applying $\psi^{k}$ to the walk $q\smash{\stackrel{{\scriptstyle\text{\tiny{$u$}}}}{{\relbar\joinrel\longrightarrow}}}\psi(q)\subseteq\Gamma(q,x_{0})$, we deduce that $q\smash{\stackrel{{\scriptstyle\text{\tiny{$u^{\omega}$}}}}{{\relbar\joinrel\longrightarrow}}}\subseteq\Gamma(q,x_{0})$ is an infinite walk that is not a circuit. ∎
To state the next lemma we need to introduce some definitions. For any end-cone type $\Gamma_{i}$ that is end-isomorphic to some end-cone $\Gamma(v,x_{0})$ with $\|v\|_{x_{0}}=n$ there are $k_{i}\leq 2|A|$ second level end-cones $\Gamma(q_{1},x_{0}),\ldots,\Gamma(q_{k_{i}},x_{0})$ contained in $\Gamma(v,x_{0})$ with $\|q_{j}\|_{x_{0}}=n+1$. Roughly speaking we are taking the end-cones of $\Gamma_{i}$ whose frontier vertices are at distance one from $\Delta_{i}$. Put $\gamma=\max\{k_{i},i\in[1,N]\}$ and $\delta=\max\{|\Delta_{i}|,i\in[1,N]\}$. A subgraph $\Lambda$ of $\Gamma$ is called $g$-periodic if $\Lambda$ is equal to the subgraph generated by the circuit $\pi=y\smash{\stackrel{{\scriptstyle\text{\tiny{$g^{k}$}}}}{{\relbar\joinrel\longrightarrow}}}y$ for some integer $k\geq 1$. The order $o(g,\Lambda)$ is the smallest integer $m$ such that $\Lambda$ is equal to the subgraph generated by the vertices and edges of the circuit $\pi=y\smash{\stackrel{{\scriptstyle\text{\tiny{$g^{m}$}}}}{{\relbar\joinrel\longrightarrow}}}y$. With a slight abuse of notation, from now on, we identify the subgraph $\Lambda$ with this minimal circuit $\pi$. We have the following lemma.
@@
For any $g$-periodic subgraph $\pi$ of a context-free graph $\Gamma$ we have $o(g,\pi)\leq|g|\delta\gamma^{N|g|}$.
@@
Note that if we cyclically shift $g$ into a word $g^{\prime}$, then $\pi$ is also a $g^{\prime}$-periodic subgraph with $o(g^{\prime},\pi)=o(g,\pi)$. Thus, by a suitable cyclic shift of $g$, we may assume without loss of generality that the subgraph
$\pi=y_{1}\smash{\stackrel{{\scriptstyle\text{\tiny{$g^{k}$}}}}{{\relbar\joinrel\longrightarrow}}}y_{1}=y_{1}\smash{\stackrel{{\scriptstyle\text{\tiny{$g$}}}}{{\relbar\joinrel\longrightarrow}}}\ldots\smash{\stackrel{{\scriptstyle\text{\tiny{$g$}}}}{{\relbar\joinrel\longrightarrow}}}y_{i}\smash{\stackrel{{\scriptstyle\text{\tiny{$g$}}}}{{\relbar\joinrel\longrightarrow}}}\ldots y_{k}\smash{\stackrel{{\scriptstyle\text{\tiny{$g$}}}}{{\relbar\joinrel\longrightarrow}}}y_{1}$
with $k=o(g,\pi)$, is contained in the end-cone $\Gamma(y_{1},x_{0})$. Suppose by contradiction that $o(g,\pi)>|g|\delta\gamma^{N|g|}$. We first observe that at most $|g|$ vertices of $\pi$ may be equal to the same vertex $v\in\Gamma$. This is true for $y_{1}$, since $\pi$ is defined to be the minimal circuit. To see this for a generic vertex of $\pi$, we say that a vertex $y_{j}\in\pi$, is of type $g^{\prime}$, where $g^{\prime}$ is a proper prefix of $g$, if there is a sub-walk
@@
$y_{1}\smash{\stackrel{{\scriptstyle\text{\tiny{$g^{t}$}}}}{{\relbar\joinrel\longrightarrow}}}y_{t}\smash{\stackrel{{\scriptstyle\text{\tiny{$g^{\prime}$}}}}{{\relbar\joinrel\longrightarrow}}}y_{j}\subseteq\pi$
for some integer $t\geq 0$. Now, if there are more than $|g|$ vertices of $\pi$ that are equal to the same vertex $v\neq y_{1}$, there would be two vertices, say $y_{i},y_{j}$, $j>i$ such that
@@
$\pi=y_{1}\smash{\stackrel{{\scriptstyle\text{\tiny{$g^{s}g^{\prime}$}}}}{{\relbar\joinrel\longrightarrow}}}y_{i}\smash{\mathop{\relbar\joinrel\relbar\joinrel\longrightarrow}\limits^{g^{\prime\prime}g^{\ell}g^{\prime}}}y_{j}\smash{\stackrel{{\scriptstyle\text{\tiny{$g^{\prime\prime}g^{t}$}}}}{{\relbar\joinrel\longrightarrow}}}y_{1}$
with $g^{\prime}g^{\prime\prime}=g$ and $y_{i}=y_{j}=v$. However, the shorter circuit $y_{i}\smash{\mathop{\relbar\joinrel\relbar\joinrel\longrightarrow}\limits^{g^{\prime\prime}g^{\ell}g^{\prime}}}y_{j}$ contradicts the minimality of $k$ since $y_{1}\smash{\stackrel{{\scriptstyle\text{\tiny{$g^{\ell+1}$}}}}{{\relbar\joinrel\longrightarrow}}}y_{1}$ would be a circuit with $\ell+1<k$. Let us now estimate the parameter
$\lambda=\max\{\|y_{j}\|_{x_{0}}-\|y_{1}\|_{x_{0}}:y_{j}\in\pi\}$
By the previous remark each vertex in $\pi$ passes from a vertex of $\Gamma$ at most $|g|$ times. Moreover, the possible number of distinct vertices through which the circuit $\pi$ passes is at most $\delta\gamma^{\lambda}$ since there are at most $\delta$ vertices in $\Delta(v,x_{0})$ and each time we are in a vertex in a frontier set, we may choose to move to the second level end-cones of which there are at most $\gamma$. In conclusion, we have that the number of vertices $k$ listed in $\pi$ are at most
$k\leq|g|\delta\gamma^{\lambda}$
which together with the inequality $k>|g|\delta\gamma^{N|g|}$ yields to $\lambda>N|g|$. Now, by the fact that $\Gamma(y_{1},x_{0})$ is quasi isometric to a tree, we have that any sub-walk of $\pi$ starting at some $y_{j}\in\Delta(y_{j},x_{0})$ and ending at $y_{1}$ must pass through some vertex $y_{s}\in\Delta(y_{j},x_{0})$ that belongs to the same set of frontier vertices of $y_{j}$. A noose of $\pi$ at $y_{j}$ is the sub-walk $\Pi(y_{j})$ of $\pi$ starting at $y_{j}$ and ending at $y_{s}\in\Delta(y_{j},x_{0})$ with $s$ minimum. It is clear that $\Pi(y_{j})\subseteq\Gamma(y_{j},x_{0})$. Fix a vertex $\overline{y}$ of $\pi$ with $\|\overline{y}\|_{x_{0}}-\|y_{1}\|_{x_{0}}=\lambda$. We say that a noose $\Pi(y_{j})$ is marked if it contains $\overline{y}$. It is clear that there is a vertex $z_{1}\in\Delta(y_{1},x_{0})$ such that $\Pi(z_{1})$ is a marked noose. There is a marked noose $\Pi(z_{2})$ with $\Pi(z_{2})\subseteq\Pi(z_{1})$ and $\|z_{2}\|_{x_{0}}=\|z_{1}\|_{x_{0}}+1$. Indeed, every sub-walk ending at $\overline{y}$ starting at $z_{1}$ must pass from the frontier vertices $z_{2}$ of a second level end-cone, and if every noose contained in $\Pi(z_{1})$ were not be marked, then $\overline{y}$ would not be contained in $\Pi(z_{1})$, a contradiction. Continuing in this way, we may find a sequence of marked nooses:
$\Pi(z_{1})\supseteq\Pi(z_{2})\supseteq\ldots\supseteq\Pi(z_{\lambda-1})\supseteq\Pi(z_{\lambda})$
where the last one may be an empty walk unless $\Pi(z_{\lambda})$ is of the form $z_{\lambda}\smash{\stackrel{{\scriptstyle\text{\tiny{$a$}}}}{{\relbar\joinrel\longrightarrow}}}z^{\prime}$ for some $z^{\prime}\in\Delta(z_{\lambda},x_{0})$. Now, let us color each $\Pi(z_{i})$ with the pair $(\Gamma_{s},g^{\prime})$ whenever $z_{i}$ is of type the proper prefix $g^{\prime}$ of $g$ and $\Gamma_{s}$ is the end-cone type with $\Gamma_{s}\simeq_{end}\Gamma(z_{i},x_{0})$. Now, since we have shown before that $\lambda>N|g|$ which is the number of possible colors, we conclude that there are two marked nooses, say $\Pi(z_{i}),\Pi(z_{j})$, $i<j$, colored with the same color, say $(\Gamma_{t},g^{\prime})$. Since $\Pi(z_{j})\subseteq\Pi(z_{i})$, there is a sub-walk
$z_{i}\smash{\stackrel{{\scriptstyle\text{\tiny{$g^{\prime\prime}g^{h}g^{\prime}$}}}}{{\relbar\joinrel\longrightarrow}}}z_{j}\subseteq\Pi(z_{i})\subseteq\Gamma(z_{i},x_{0})$
with $g^{\prime}g^{\prime\prime}=g$. Therefore, since $\Gamma(z_{i},x_{0}),\Gamma(z_{j},x_{0})$ are two self-similar end-cones of $\Gamma$, by Lemma 5.1 there is an infinite walk
$z_{i}\smash{\mathop{\relbar\joinrel\relbar\joinrel\relbar\joinrel\longrightarrow}\limits^{(g^{\prime\prime}g^{h}g^{\prime})^{\omega}}}$
with $z_{i}\in\pi$ that is not a circuit, hence $\pi$ is also $g^{\prime\prime}g^{\prime}$-periodic and $\pi=z_{i}\smash{\mathop{\relbar\joinrel\relbar\joinrel\relbar\joinrel\longrightarrow}\limits^{(g^{\prime\prime}g^{\prime})^{\omega}}}$, a contradiction. Thus, the bound $o(g,\pi)\leq|g|\delta\gamma^{N|g|}$ holds. ∎
The following proposition gives some information regarding torsion elements.
@@
  • let $\Gamma(p,x_{0}),\Gamma(q,x_{0})$ be two distinct end-isomorphic self-similar end-cones of $\Gamma$, and let $q\smash{\stackrel{{\scriptstyle\text{\tiny{$u$}}}}{{\relbar\joinrel\longrightarrow}}}p\subseteq\Gamma(q,x_{0})\subseteq$ be a walk. Then, the element $u\in\mathcal{G}(\Gamma)$ has infinite order.
  • Let $g\in\mathcal{G}(\Gamma)$ be an element of finite order, then the order $o(g)$ of the element $g$ satisfies $\frac{o(g)}{\log_{2}(o(k))}<|g|\delta\gamma^{N|g|}$.
@@
The first fact follows from Lemma 5.1 since $q\smash{\stackrel{{\scriptstyle\text{\tiny{$u^{\omega}$}}}}{{\relbar\joinrel\longrightarrow}}}$ is an infinite walk that is not a circuit we have $u^{i}\cdot q\neq u^{j}\cdot q$ for any $i,j$ with $i\neq j$. Let us prove the bound on the torsion element $g$. Let $o(g)=p_{1}^{\alpha_{1}}\ldots p_{s}^{\alpha_{s}}$ be the prime factorization of $o(g)$. We prove that for any $p_{i}^{\alpha_{i}}$ appearing in the factorization of $o(g)$ we have $p_{i}^{\alpha_{i}}\leq|g|\delta\gamma^{N|g|}$. By definition of the group $\mathcal{G}(\Gamma)$, there is a $g$-periodic subgraph $\pi$ with $o(g,\pi)=m$ and such that $m$ is a multiple of $p_{i}^{\alpha_{i}}$. Thus, by Lemma 5.2 we have
$p_{i}^{\alpha_{i}}\leq m\leq o(g,\pi)\leq|g|\delta\gamma^{N|g|}$
The number of primes appearing in the prime factorization on a positive integer $n$ is maximized by the integers of the form $n=2\cdot 3\cdot\ldots\cdot p_{k}$, thus by a rough estimation $n$ has at most $\log_{2}(n)$ primes appearing in its prime factorization. Hence, the upper bound of the statement follows. ∎

### v3

@@ 5. Detecting torsion
We give some structural results focusing on the existence of torsion elements and on bounds on the order of such elements. Henceforth, we consider a context-free $A$-graph $\Gamma$ with $N$ end-cone types $\Gamma_{1},\ldots,\Gamma_{N}$ with the corresponding frontier vertices $\Delta_{1},\ldots,\Delta_{N}$ with $\Delta_{i}\subseteq\Gamma_{i}$. For any end-cone type $\Gamma_{i}$ that is end-isomorphic to some end-cone $\Gamma(v,x_{0})$ with $\|v\|_{x_{0}}=n$ there are $k_{i}\leq 2|A|$ second level end-cones $\Gamma(q_{1},x_{0}),\ldots,\Gamma(q_{k_{i}},x_{0})$ contained in $\Gamma(v,x_{0})$ with $\|q_{j}\|_{x_{0}}=n+1$. Roughly speaking we are taking the end-cones of $\Gamma_{i}$ whose frontier vertices are at distance one from $\Delta_{i}$. Put $\gamma=\max\{k_{i},i\in[1,N]\}$ and $\delta=\max\{|\Delta_{i}|,i\in[1,N]\}$.
We say that two end-cones $\Gamma(p,x_{0}),\Gamma(q,x_{0})$ are nested if $\Gamma(p,x_{0})\subseteq\Gamma(q,x_{0})$ and there is an end-isomorphism $\psi:\Gamma(q,x_{0})\to\Gamma(p,x_{0})$ with $\psi(q)=p$. We have the following lemma.
@@
Let $\Gamma(p,x_{0}),\Gamma(q,x_{0})$ be two distinct end-isomorphic nested end-cones of $\Gamma$. Then, for any walk $q\smash{\stackrel{{\scriptstyle\raisebox{-2.71246pt}{\tiny$u$}}}{{\relbar\joinrel\longrightarrow}}}p\subseteq\Gamma(q,x_{0})$ there is an infinite walk $q\smash{\stackrel{{\scriptstyle\raisebox{-2.71246pt}{\tiny$u^{\omega}$}}}{{\relbar\joinrel\longrightarrow}}}\ldots\subseteq\Gamma(q,x_{0})$ that is not a circuit.
@@
Indeed, let $\psi:\Gamma(q,x_{0})\to\Gamma(p,x_{0})$ be the end-isomorphism with $\psi(q)=p$. Consider a geodesic $q\smash{\stackrel{{\scriptstyle\raisebox{-2.71246pt}{\tiny$g$}}}{{\relbar\joinrel\longrightarrow}}}\psi(q)\subseteq\Gamma(q,x_{0})$ connecting $q$ with $\psi(q)=p$. We claim that also $q\smash{\stackrel{{\scriptstyle\raisebox{-2.71246pt}{\tiny$g$}}}{{\relbar\joinrel\longrightarrow}}}\psi(q)\smash{\stackrel{{\scriptstyle\raisebox{-2.71246pt}{\tiny$g$}}}{{\relbar\joinrel\longrightarrow}}}\psi^{2}(q)$ is a geodesic. Any geodesic $q\smash{\stackrel{{\scriptstyle\raisebox{-2.71246pt}{\tiny$u$}}}{{\relbar\joinrel\longrightarrow}}}\psi^{2}(q)\subseteq\Gamma(q,x_{0})$ has to cross a vertex $v\in\Delta(\psi(q),x_{0})$, thus the previous geodesic factors as $q\smash{\stackrel{{\scriptstyle\raisebox{-2.71246pt}{\tiny$u_{1}$}}}{{\relbar\joinrel\longrightarrow}}}v\smash{\stackrel{{\scriptstyle\raisebox{-2.71246pt}{\tiny$u_{2}$}}}{{\relbar\joinrel\longrightarrow}}}\psi^{2}(q)$ with $q\smash{\stackrel{{\scriptstyle\raisebox{-2.71246pt}{\tiny$u_{1}$}}}{{\relbar\joinrel\longrightarrow}}}v\subseteq\Gamma(q,x_{0})$ and $v\smash{\stackrel{{\scriptstyle\raisebox{-2.71246pt}{\tiny$u_{2}$}}}{{\relbar\joinrel\longrightarrow}}}\psi^{2}(q)\subseteq\Gamma(\psi(q),x_{0})$. Since both $q\smash{\stackrel{{\scriptstyle\raisebox{-2.71246pt}{\tiny$g$}}}{{\relbar\joinrel\longrightarrow}}}\psi(q)$ and $\psi(q)\smash{\stackrel{{\scriptstyle\raisebox{-2.71246pt}{\tiny$g$}}}{{\relbar\joinrel\longrightarrow}}}\psi^{2}(q)$ are geodesics, respectively in $\Gamma(q,x_{0})$ and $\Gamma(\psi(q),x_{0})$ we deduce $|u_{1}|=|g|$ and $|u_{2}|=|g|$. Hence, $q\smash{\stackrel{{\scriptstyle\raisebox{-2.71246pt}{\tiny$g$}}}{{\relbar\joinrel\longrightarrow}}}\psi(q)\smash{\stackrel{{\scriptstyle\raisebox{-2.71246pt}{\tiny$g$}}}{{\relbar\joinrel\longrightarrow}}}\psi^{2}(q)$ is also a geodesic. The same argument can be extended to show that $q\smash{\stackrel{{\scriptstyle\raisebox{-2.71246pt}{\tiny$g^{k}$}}}{{\relbar\joinrel\longrightarrow}}}\psi^{k}(q)\subseteq\Gamma(q,x_{0})$ is a geodesic for any $k\geq 1$. Now, by applying $\psi^{k}$ to the walk $q\smash{\stackrel{{\scriptstyle\raisebox{-2.71246pt}{\tiny$u$}}}{{\relbar\joinrel\longrightarrow}}}\psi(q)\subseteq\Gamma(q,x_{0})$, we deduce that $q\smash{\stackrel{{\scriptstyle\raisebox{-2.71246pt}{\tiny$u^{\omega}$}}}{{\relbar\joinrel\longrightarrow}}}\subseteq\Gamma(q,x_{0})$ is an infinite walk that is not a circuit. ∎
A subgraph $\Lambda$ of $\Gamma$ is called $g$-periodic if $\Lambda$ is equal to the subgraph induced by the circuit $\pi=y\smash{\stackrel{{\scriptstyle\raisebox{-2.71246pt}{\tiny$g^{k}$}}}{{\relbar\joinrel\longrightarrow}}}y$ for some integer $k\geq 1$. The order $o(g,\Lambda)$ is the smallest integer $m$ such that $\Lambda$ is equal to the subgraph induced by the vertices and edges of the circuit $\pi=y\smash{\stackrel{{\scriptstyle\raisebox{-2.71246pt}{\tiny$g^{m}$}}}{{\relbar\joinrel\longrightarrow}}}y$. A clip is a walk $y_{1}\smash{\stackrel{{\scriptstyle\raisebox{-2.71246pt}{\tiny$u$}}}{{\relbar\joinrel\longrightarrow}}}y_{2}\subseteq\Gamma(y_{1},x_{0})$ with $y_{1},y_{2}\in\Delta(y_{1},x_{0})$. For a clip $p$, we denote by $(\chi(p),\subseteq)$ the poset of all subwalks of $p$ which are clips ordered by inclusion.
@@
For a clip $p=y_{1}\smash{\stackrel{{\scriptstyle\raisebox{-2.71246pt}{\tiny$u$}}}{{\relbar\joinrel\longrightarrow}}}y_{2}$, let $h(\chi(p))$ be the length of a maximal chain in $\chi(p)$. Then, $h(\chi(p))\geq\max\{d(y_{1},x):x\in p\}+1$.
@@
Let $k=\max\{d(y_{1},x):x\in p\}$. We proceed by induction on $k$. The case $k=0$ is trivial. Assume that $k>0$. Let $q_{1}$ and $q_{2}$ denote, respectively, the first and the last vertices of $p$ such that $d(y_{1},q_{1})=d(y_{1},q_{2})=k-1$. Such vertices exist because, by the tree-like structure of $\Gamma$, the walk $p$ must pass through the set of frontier vertices at distance k1k-1 both when entering and when exiting $\Gamma(q_{1},x_{0})$ which is a second level end-cone of $\Gamma(y_{1},x_{0})$. Then the subwalk $p^{\prime}=q_{1}\smash{\stackrel{{\scriptstyle\raisebox{-2.71246pt}{\tiny$u^{\prime}$}}}{{\relbar\joinrel\longrightarrow}}}q_{2}$ of $p$ is a clip with $\max\{d(q_{1},x):x\in p^{\prime}\}=k-1$. By the induction hypothesis, we have $h(\chi(p^{\prime}))\geq k$. Hence, since $p^{\prime}\subsetneq p$, it follows that $h(\chi(p))\geq h(\chi(p^{\prime}))+1\geq k+1$. ∎
Lemma 5.3.
For any $g$-periodic subgraph $\pi$ of a context-free graph $\Gamma$ we have $o(g,\pi)<\left(\frac{|g|-1}{|g|}\delta\right)\gamma^{\delta^{2}N|g|}$.
Proof.
Let $\pi$ be the subgraph induced by the walk $p=y\smash{\stackrel{{\scriptstyle\raisebox{-2.71246pt}{\tiny$g^{\ell}$}}}{{\relbar\joinrel\longrightarrow}}}y$, where $\ell=o(g,\pi)$. Note that, without loss of generality, by cyclically shifting $g$ if necessary, we may assume that $p\subseteq\Gamma(y,x_{0})$, so that $p$ is a clip. For any clip $\lambda\in\chi(p)$ with $\lambda=y_{1}\smash{\stackrel{{\scriptstyle\raisebox{-2.71246pt}{\tiny$u$}}}{{\relbar\joinrel\longrightarrow}}}y_{2}$ and $y_{1},y_{2}\in\Delta(y_{1},x_{0})$, we say that $\lambda$ is colored by $(x_{1},x_{2},\Gamma_{i},g^{\prime})$ if the following conditions hold:
  • $\lambda\subseteq\Gamma(y_{1},x_{0})$ and there exists an end-isomorphism $\psi:\Gamma(y_{1},x_{0})\to\Gamma_{i}$ such that $x_{1}=\psi(y_{1}),x_{2}=\psi(y_{2})\in\Delta_{i}$;
@@
  • $g^{\prime}$ is a prefix of $g$, and there exists an integer $s\geq 0$ such that $y\smash{\stackrel{{\scriptstyle\raisebox{-2.71246pt}{\tiny$g^{s}g^{\prime}$}}}{{\relbar\joinrel\longrightarrow}}}y_{1}$.
Let us prove first the following claim:
(2) $\mbox{if }\lambda_{1},\lambda_{2}\in\chi(p),\lambda_{1}\subsetneq\lambda_{2},\mbox{then }\lambda_{1},\lambda_{2}\mbox{ have different colors.}$
Suppose by contradiction that claim (2) does not hold, and $\lambda_{1},\lambda_{2}$ have the same color $(x_{1}.x_{2},\Gamma_{i},g^{\prime})$. Therefore, $\lambda_{1}$ and $\lambda_{2}$ can be factorized as follows:
@@
$\lambda_{2}=y_{1}\smash{\stackrel{{\scriptstyle\raisebox{-2.71246pt}{\tiny$g^{\prime\prime}$}}}{{\relbar\joinrel\longrightarrow}}}q_{1}\smash{\stackrel{{\scriptstyle\raisebox{-2.71246pt}{\tiny$g^{m}$}}}{{\relbar\joinrel\longrightarrow}}}q_{2}\smash{\stackrel{{\scriptstyle\raisebox{-2.71246pt}{\tiny$g^{\prime}$}}}{{\relbar\joinrel\longrightarrow}}}y_{1}^{\prime}\smash{\stackrel{{\scriptstyle\raisebox{-2.71246pt}{\tiny$g^{\prime\prime}u_{2}$}}}{{\relbar\joinrel\longrightarrow}}}y_{2}^{\prime}\smash{\stackrel{{\scriptstyle\raisebox{-2.71246pt}{\tiny$u_{3}$}}}{{\relbar\joinrel\longrightarrow}}}y_{2},\qquad\lambda_{1}=y_{1}^{\prime}\smash{\stackrel{{\scriptstyle\raisebox{-2.71246pt}{\tiny$g^{\prime\prime}u_{2}$}}}{{\relbar\joinrel\longrightarrow}}}y_{2}^{\prime},$
for some $m\geq 0$, where $g^{\prime}g^{\prime\prime}=g$. Since $\lambda_{1}$ and $\lambda_{2}$ have the same color, there exists an end-isomorphism $\psi:\Gamma(y_{1},x_{0})\to\Gamma(y_{1}^{\prime},x_{0})$ between two nested end-cones $\Gamma(y_{1}^{\prime},x_{0})\subseteq\Gamma(y_{1},x_{0})$. Hence, by Lemma 5.1, there exists an infinite walk $y_{1}\smash{\mathop{\relbar\joinrel\relbar\joinrel\relbar\joinrel\longrightarrow}\limits^{\raisebox{-2.1097pt}{\tiny$(g^{\prime\prime}g^{m}g^{\prime})^{\omega}$}}}\ldots=y_{1}\smash{\mathop{\relbar\joinrel\relbar\joinrel\relbar\joinrel\longrightarrow}\limits^{\raisebox{-2.1097pt}{\tiny$(g^{\prime\prime}g^{\prime})^{\omega}$}}}\ldots$ that is not a circuit, and this implies that $\pi$ is infinite, a contradiction.
By the determinism and the minimality of $\ell$ each vertex $q\in V(\pi)$ occurs at most |g|1|g|-1 times in the vertices appearing in the description of the walk $p=y\smash{\stackrel{{\scriptstyle\raisebox{-2.71246pt}{\tiny$g^{\ell}$}}}{{\relbar\joinrel\longrightarrow}}}y$. Hence, the number of vertices appearing in $p$ is at most $(|g|-1)\,|V(\pi)|$. We claim that $|V(\pi)|\leq\delta\gamma^{C(g)}$, where $C(g)$ denotes the number of possible colors. Suppose, for the sake of contradiction, that $|V(\pi)|>\delta\gamma^{C(g)}$. Since the number of vertices in $\Gamma(y,x_{0})$ at distance $n$ satisfies $|\Gamma(y)\cap D_{n}(y)|<\delta\gamma^{n}$, there must exist a vertex $v\in p$ with $d(v,y)\geq C(g)$. Then, by Lemma 5.2, we conclude that the length of a maximal chain $\lambda_{1}\subseteq\lambda_{2}\subseteq\ldots\subseteq\lambda_{m}$ in $\chi(p)$ is strictly greater than $C(g)+1$, that is, $m>C(g)$. Therefore, by the pigeonhole principle, there exist $\lambda_{i}\subsetneq\lambda_{j}$ with the same color, contradicting our claim (2). The lemma now follows by noting that $C(g)\leq\delta^{2}N|g|$ and the inequality $|g|\ell\leq(|g|-1)\,|V(\pi)|$
The following proposition provides information about the order of a element.
@@
  • let $\Gamma(p,x_{0}),\Gamma(q,x_{0})$ be two distinct end-isomorphic nested end-cones of $\Gamma$, and let $q\smash{\stackrel{{\scriptstyle\raisebox{-2.71246pt}{\tiny$u$}}}{{\relbar\joinrel\longrightarrow}}}p\subseteq\Gamma(q,x_{0})\subseteq$ be a walk. Then, the element $u\in\mathcal{G}(\Gamma)$ has infinite order.
  • Let $g\in\mathcal{G}(\Gamma)$ be an element of finite order, then the order $o(g)$ of the element $g$ satisfies $o(g)<e^{-N\delta}\delta^{N\delta|g|}\gamma^{\delta^{3}N^{2}|g|^{2}}=O(exp(|g|^{2}))$.
@@
The first claim follows from Lemma 5.1: since $q\smash{\stackrel{{\scriptstyle\raisebox{-2.71246pt}{\tiny$u^{\omega}$}}}{{\relbar\joinrel\longrightarrow}}}\ldots$ is an infinite walk that is not a circuit, we have $u^{i}\cdot q\neq u^{j}\cdot q$ for all $i\neq j$. Let us now prove the bound on the torsion element $g$. Note that $o(g)$ is upper bounded by the least common multiple of all possible values $o(g,\pi)$, where $\pi$ ranges over all $g$-periodic subgraphs. By Lemma 5.3, these subgraphs are, up to end-isomorphism, finitely many; denote this finite family by $C=\{\pi_{1},\ldots,\pi_{k}\}$. Each $\pi\in C$ is entirely contained in some end-cone $\Gamma_{i}$ and is the subgraph induced by the walk $y\smash{\stackrel{{\scriptstyle\raisebox{-2.71246pt}{\tiny$g_{1}^{\ell}$}}}{{\relbar\joinrel\longrightarrow}}}y$, where $g_{1}$ is a cyclic shift of $g$. Since there are at most $|g|$ cyclic shifts of $g$, there can be at most $|g|$ $g$-periodic graphs $\pi\subseteq\Gamma_{i}$ with $y\in\Delta_{i}\cap V(\pi)$. Therefore, we may conclude that the cardinality $|C|\leq|g|N\delta$. Hence, the least common multiple of all possible values of $o(g,\pi)$, $\pi\in C$, can be upper bounded by $(\frac{|g|-1}{|g|}\delta\gamma^{\delta^{2}N|g|})^{|g|N\delta}$. Thus, we conclude $o(g)<(\frac{|g|-1}{|g|}\delta\gamma^{\delta^{2}N|g|})^{|g|N\delta}\leq e^{-N\delta}\delta^{N\delta|g|}\gamma^{\delta^{3}N^{2}|g|^{2}}$. ∎